OpenStudy (anonymous):

PLEASE CHECK ! MEDAL WILL BE GIVEN !! Solve, using the substitution method. y = 4x + 6 5x – y = 6 There are an infinite number of solutions. There is no solution. (18, 12) (12, 54) ANSWER- A ? ?

jimthompson5910 (jim_thompson5910):

5x – y = 6 5x – ( y ) = 6 5x – ( 4x + 6 ) = 6 ... plug in y = 4x + 6 (ie replace y with 4x+6) 5x - 4x - 6 = 6 ... ... ... x = ???

OpenStudy (anonymous):

-6 ?

jimthompson5910 (jim_thompson5910):

what is 5x - 4x equal to?

OpenStudy (anonymous):

1x ? lol

jimthompson5910 (jim_thompson5910):

or just x, good

jimthompson5910 (jim_thompson5910):

so we have x - 6 = 6 now

jimthompson5910 (jim_thompson5910):

what's the next step?

OpenStudy (anonymous):

multiply both sides by 6 ?

jimthompson5910 (jim_thompson5910):

what operation undoes subtraction?

OpenStudy (anonymous):

i really dont know and im runing out of time is there a chance maybe if i can just get an answer? please ? ?

jimthompson5910 (jim_thompson5910):

you have to add 6 to both sides to undo the -6 what do you get when you do that?

OpenStudy (anonymous):

12

jimthompson5910 (jim_thompson5910):

so x = 12 now plug this into y = 4x+6 to get y = 4x + 6 y = 4*12+6 y = 48+6 y = 54

jimthompson5910 (jim_thompson5910):

therefore, x = 12 and y = 54

OpenStudy (anonymous):

thanks you so much

jimthompson5910 (jim_thompson5910):

you're welcome